# A Survey of Generative AI for Intelligent Transportation Systems

### v1

@@ 2.1.4. Energy-Based Models
EBMs (LeCun et al., 2006; Du and Mordatch, 2019) focus on defining an energy function that assigns a scalar value to each data point. The energy function measures how well a data point fits the model, with lower energy values indicating a better fit. EBMs aim to assign lower energy values to realistic data points and higher energy values to data points that are not part of the true distribution. In other words, they can approximate the likelihood of the observed data and offer a quantitative assessment of how well a given data point aligns with the learned distribution.
For the observed data $\mathbf{x}$, its probability density function $p_{\theta}(\mathbf{x})$ is defined via the Boltzmann distribution:
(4)4( 4 ) $p_{\theta}(\mathbf{x})=\frac{e^{-E_{\theta}(\boldsymbol{\mathbf{x}})}}{\int_{\tilde{\mathbf{x}}\in\mathcal{X}}e^{-E_{\theta}(\tilde{\mathbf{x}})}},$
where $E_{\theta}(\mathbf{x})$ is an energy function with parameters $\theta$.
A common approach to optimize EBMs is contrastive divergence. In contrastive divergence, the model first samples a sample from the positive data distribution and another sample from the negative data distribution. Subsequently, the model updates its parameters by comparing the energy values of these two samples. The objective is to decrease the energy value of positive samples while increasing the energy value of negative samples. This process is typically repeated iteratively to gradually optimize the model.
@@ 2.1.5. Diffusion Probabilistic Model
The essence of DPMs (Sohl-Dickstein et al., 2015) lies in its step-by-step denoising process that gradually recovers the underlying original data samples.
[…]
In the forward process, a Markov chain successively adds Gaussian noise to the data $\mathbf{x}_{0}$. Hence, we can obtain the approximate posterior $q\left(\boldsymbol{x}_{1:T}\mid\boldsymbol{x}_{0}\right)$:
(5)5( 5 ) $q\left(\boldsymbol{x}_{1:T}\mid\boldsymbol{x}_{0}\right)=\prod_{t=1}^{T}q\left(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t-1}\right),$
where $q\left(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t-1}\right)$ is a Gaussian distribution.
In the reverse process, the original data point is recovered by iteratively removing noise from the noisy data. Specifically, this process is structured as a Markov chain, leveraging learnable Gaussian transitions that initiate from the distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$:
(6)6( 6 ) $p_{\theta}\left(\boldsymbol{x}_{0:T}\right)=p\left(\boldsymbol{x}_{T}\right)\prod_{t=T}^{1}p_{\theta}\left(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t}\right),$
where $p_{\theta}(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t})$ represents a reverse process.
During each training iteration, the network attempts to predict the exact noise added to the data at a specific diffusion step. The training objective is to minimize the difference between the network’s predicted noise and the actual noise.
@@ 2.1.6. Generative Models from Physical Process
GenPhys (Liu et al., 2023c) transform physical partial differential equations (PDEs) into generative models. They provide a more general generative framework, with diffusion probabilistic model being a special case within the GenPhys framework.
Given observed data sampled from the probability distribution $p_{data}(\mathbf{x})$, the task aims to produce new sample from $p_{data}(\mathbf{x})$. A continuous physical process can be described by
(7)7( 7 ) $\frac{d\mathbf{x}}{dt}=\mathbf{v}(\mathbf{x},t),$
where $\mathbf{v}$ represents a velocity field.
This process dynamically evolves the probability distribution $p(\mathbf{x},t)$ according to the following equation:
(8)8( 8 ) $\frac{\partial p(\mathbf{x},t)}{\partial t}+\nabla\cdot[p(\mathbf{x},t)\mathbf{v}(\mathbf{x},t)]-R(\mathbf{x},t)=0,$
where $R(\mathbf{x},t)$ can be interpreted as birth or death. When $R>0$, it means indicates the birth of particles in the forward process and their subsequent death in the backward process, and vice versa. This equation is known as density flow equation. The goal is to design $p(\mathbf{x},t),\mathbf{v}(\mathbf{x},t)$ and $R(\mathbf{x},t)$ in a manner that fulfills two conditions. First, it should ensure that $p(\mathbf{x},0)=p_{data}(\mathbf{x})$. The second condition is that $p_{\text{prior}}(\mathbf{x})\equiv p(\mathbf{x},T)$ becomes asymptotically independent of $p_{\text{data}}(\mathbf{x})$ as $T\rightarrow\infty$.
Continuous physical processes are governed by PDEs, which is written as:
(9)9( 9 ) $\hat{L}\phi\equiv F\left(\phi,\phi_{t},\phi_{tt},\nabla\phi,\nabla^{2}\phi,\ldots\right)=f(\mathbf{x},t),$
where $\hat{L}$ represents a differential operator on the scalar function $\phi(\mathbf{x},t)$. $f(\mathbf{x},t)$ is the source term.
Next, the crucial step is to map the solutions of physical PDEs to those of density flows. This requires density flows and physical PDEs are equivalent equations. The final step is to produce $\mathbf{x}(0)$ by the reverse physical process. To begin, $\mathbf{x}(T)$ is drawn from the distribution $p(\mathbf{x},T)$. Then, we simulate the backward evolution of $\frac{d\mathbf{x}}{dt}=\mathbf{s}_{\theta}(\mathbf{x},t)$ from $t=T$ to $t=0$ using the branching process $W_{\alpha}(\mathbf{x},t)$. This simulation ultimately produces $\mathbf{x}(0)$.
@@ 2.1.7. Generative Pre-trained Transformer
Generative Pre-trained Transformer (GPT) is a neural network framework based on Transformer, specifically developed for tasks involving natural language processing (NLP) and generation. GPT models are pre-trained on a large number of textual datasets and are capable of generating text that is both coherent and contextually meaningful. GPT utilizes a transformer architecture (Vaswani et al., 2017) with decoder layers. Each decoder layer consists of self-attention mechanisms and feedforward neural networks. These mechanisms allow the model to capture long-range dependencies in sequential data like text. To handle sequential input data, GPT includes positional encodings and learned word embeddings. Layer normalization is applied after each sub-layer in the transformer decoder to stabilize training. GPT employs an autoregressive generation approach, producing output tokens sequentially, building upon previously produced tokens. An important feature of GPT is parameter sharing across its decoder layers, ensuring consistency in data representation. GPT models undergo unsupervised pre-training on extensive text corpora and can be fine-tuned for specific tasks with labeled data. Their scalability has led to increasingly large models, such as GPT-2 (Nguyen et al., 2021), GPT-3 (Brown et al., 2020) and GPT-4 (OpenAI, 2023), achieving superior performance across various NLP tasks.
@@ 2.1.8. Comparative Analysis of Generative AI Techniques
In previous subsections, we have introduced several major generative AI techniques. Each of them has its distinct advantages and limitations. For example, VAE offer a probabilistic framework for encoding data into a lower-dimensional space but often face challenges in selecting the appropriate likelihood functions and generating high-quality images. GANs aim to generate highly realistic data samples by pitting a generator against a discriminator in a adversarial training scheme. They produce impressive results in image generation but can be challenging to train and prone to mode collapse. We summarize the advantages and disadvantages of these generative AI techniques, as shown in Table 2.
Technologies Advantages Disadvantages
Variational Autoencoder (Kingma and Welling, 2013)
- Explicit encoder and decoder
- Capable of probabilistic generation, suitable for data with uncertainty
- Can learn latent representations of data
- Training is relatively stable
- May produce blurry images
- Challenges in selecting likelihood functions
- Limited expressiveness of the prior distribution
Generative Adversarial Network (Goodfellow et al., 2020)
- Capable of generating high-quality images
- Provides diversity in the generation process
- Training process can be unstable
- Mode collapse issue is common
Normalizing Flow (Rezende and Mohamed, 2015)
- Can accurately model probability distributions
- Supports both generation and inference
- Applicable to various data types
- Training and inference complexities are high
- Requires more parameters for high-dimensional data
Energy-Based Model (LeCun et al., 2006)
- Can model complex probability distributions
- Supports both generation and inference
- Robust to outliers
- Requires careful tuning of hyperparameters
- Training process may be slow
Diffusion Probabilistic Model (Sohl-Dickstein et al., 2015)
- Robust to noise and uncertainty
- Flexible noise schedule choice
- Allows for controlled generation
- Longer generation time
- Training complex diffusion models can be slow
Generative Models from Physical Process (Liu et al., 2023c)
- Generates data based on physical processes
- Offers interpretability
- Requires prior knowledge of physical parameters
- Typically limited to specific application domains
Generative Pre-trained Transformer (Ghojogh and Ghodsi, 2020)
- Performs exceptionally well in natural language processing tasks
- Capable of generating continuous text
- Effective modeling of long-term dependencies
- Requires substantial computing resources for training
- Dependence on large training datasets
Table 2. The advantages and disadvantages of Generative AI techniques, partly referenced by the work (Zhang et al., 2023d).
@@ 2.2. Important Generation Tasks
The advancement of generative technologies has provided impressive solutions for various generative tasks. In this section, we will explore the applications of generative technologies in these generation tasks from the perspectives of text, images, videos, and cross-modal data.
@@ 2.2.1. Text Generation
Text generation focuses on using AI techniques to autonomously generate reasonable and meaningful text. This technology is often used in tasks such as machine translation, chatbots, and article summarization. In recent years, with the advancement of deep learning technologies, especially the emergence of the Transformer architecture and the GPT series of models, both the quality and diversity of text generation have significantly improved. ChatGPT (Ouyang et al., 2022), developed by OpenAI, is a specialized version of the GPT model fine-tuned for conversational interactions. Microsoft’s Bing (Adiwardana et al., 2020) has integrated ChatGPT technology, enhancing the conversational search experience for users. This integration empowers users to comprehend intricate queries, receive detailed responses, and benefits from continuous learning and optimization capabilities.
@@ 2.2.2. Image Generation
Image generation focuses on using AI algorithms to create new images or visual content using given information such as images and text. For example, generative methods leverage the power of neural networks to learn complex data representations, making them a prominent approach for restoring images to their original high-quality state (Sohl-Dickstein et al., 2015; Ledig et al., 2017; Cai and Wei, 2020; Rombach et al., 2022). Several works focus on conversion between two images, encompassing tasks such as style transfer and image morphing (Karras et al., 2019; Zhao et al., 2021b). Recently, diffusion models are utilized to address the tasks of image editing, achieving superior image generation quality (Li et al., 2022e; Ackermann and Li, 2022; Avrahami et al., 2023).
@@ 2.2.3. Video Generation
Video generation involves creating a series of frames that are semantically coherent and appear in a sequence to form a video. This task is challenging due to its requirement to deal with high-dimensional video data and simultaneously model spatial and temporal depenedencies. The related works are usually classified into two types. Unconditional video generation aims to generate new videos without requiring additional information. It achieves this by extracting valuable information from the training data. Numerous works that tackle the task of unconditional video generation utilize generative AI technology (Wang et al., 2020; Gupta et al., 2021; Wang et al., 2021a). Unlike unconditional video generation, conditional video generation incorporates additional condition to guide the process of video generation. VAE and GAN are are two popular methods used to generate videos with specific settings such as text (Li et al., 2018a), sentences (Liu et al., 2019) and captions (Pan et al., 2017).
@@ 2.2.4. Cross-modality Data Generation
Cross-modality data generation refers to generate data in one modality such as image, text and video based on the features from a different modality. The primary challenge in implementing cross-modality data generation lies in effectively modeling the relationships between different modalities, enabling control over the generation process through inputs from another modality (Foo et al., 2023). Text-to-image generation aims to generate images based on a provided textual description. Current approaches for text-to-image generation are mainly built upon GANs (Xu et al., 2018; Li et al., 2019), Transformer (Ramesh et al., 2021; Ding et al., 2021, 2022) and diffusion models (Nichol et al., 2021; Radford et al., 2021). Image-to-text generation involves generating textual description or captions that describe the content of an image. Most works adopt the encoder-decoder architectures to deal with this task. The language decoder plays a crucial role in model design. Advancements in the NLP field have led to a transition in language decoder architecture from RNNs to Transformer-based generative models (Li et al., 2020c; Zhang et al., 2021b). Text-to-video generation refers to creating videos from the textual prompts. Three commonly used approaches to address the challenge of text-to-video generation include GAN-based (Balaji et al., 2019) and VAE-based (Mittal et al., 2017), and diffusion-based (Singer et al., 2022; Rahman et al., 2023) methods.
2.3. Advantages of generative AI technology
Generative AI technology has significant advantages over traditional deep learning methods such as MLP, CNN, RNN in many aspects. In this subsection, we will discuss them in terms of data generation, creativity, data utilization, model generalization, and cross-domain applications.
Data generation. Generative AI not only has the capability to process and analyze existing data but also to generate new data samples based on the distribution of the available data. For instance, in the field of computer vision, GANs can produce highly realistic images. In natural language processing, generative pre-trained models like the GPT series are proficient at generating fluent text, which can be used for tasks such as automated content generation, chatbot systems, and translation services. Furthermore, generative AI finds applications in various domains, including audio synthesis, video generation, medical image synthesis, and simulating experiments in scientific research. While traditional deep learning methods excel in data analysis and classification tasks, they typically lack the direct capability to generate entirely new data samples. This makes generative AI particularly advantageous in data augmentation and simulation among other areas.
Creativity. Generative AI has the ability to generate novel, interesting, and diverse data, which enables it to have extensive application potential in the creative field. For example, in the field of literature, generative pre-trained models such as GPT-3 (Floridi and Chiriatti, 2020) have been able to automatically generate articles, essays, poems, and stories. Generative AI is also used in music composition (Muhamed et al., 2021; Zhang, 2020; Shih et al., 2022). Music generation models can generate various styles and types of music, from classical to popular. In the field of scientific research, generative AI is also used for new drug discovery (Cheng et al., 2021b; Bilodeau et al., 2022). By simulating molecular structures and chemical reactions, generative AI can generate potential new drug candidates, accelerating the process of drug development. Traditional deep learning methods rely more on existing data for extraction and classification, and are relatively weak in terms of creativity.
Reasoning. Generative AI methods have more advantages than traditional deep learning methods in terms of reasoning. Firstly, they often employ probabilistic modeling, which can effectively model uncertainties. Secondly, they have the ability to capture long-term dependencies. For example, the self-attention mechanism in the GPT architecture allows the model to capture information from any position in the sequence during the processing of sequence data, rather than just local dependencies. Traditional deep learning methods typically focus on end-to-end learning, with the main goal of mapping input data to output data rather than specific reasoning. Therefore, the probabilistic modeling and long-term dependency extraction of the generative AI methods make it more outstanding in reasoning.
Model generalization. Generative AI not only learns broader knowledge through large-scale data training, but also benefits from its architecture with numerous shared parameters, allowing it to transfer learned knowledge across different tasks. For example, GPT can be applied to a variety of NLP tasks including text generation, language translation, and dialogue generation, without the need to design new models for each task separately. Similarly, in computer vision, generative AI-based models can initially be trained for image classification, and then are employed for other tasks such as object detection without the requirement for redesigning network structures. Achieving such adaptability and versatility proves challenging for traditional deep learning methods, which typically require task-specific network architectures.
Cross-domain applications. Generative AI methods can enable cross-domain data generation, allowing knowledge acquired in one domain to seamlessly transfer to another. For example, Google’s Imagen uses a diffusion model to convert text descriptions into images (Ho et al., 2022). This cross-domain transfer learning provides new opportunity for knowledge sharing between different domains. In contrast, traditional deep learning methods are relatively weak in cross-domain applications. They usually require redesigning and training models specific to the new domain to adapt to the data and tasks of different domains. This leads to increased complexity and resource requirements for cross-domain knowledge transfer, limiting their extensive applicability across different domains.
@@ 3. Generative AI for Traffic Perception
Traffic perception refers to the ability of intelligent transportation systems to collect and understand various sensory information in the traffic environment. This information includes visual data from cameras, trajectory data from GPS-based devices, motion data from accelerometers, environmental data like weather and road conditions, and other traffic-related information such as the presence of other vehicles, pedestrians, and road signs. It plays a crucial role in making effective decisions and ensuring safe driving. In the context of autonomous vehicles, traffic perception typically involves the use of cameras, LiDAR, millimeter-wave radar, and other technologies to collect and process data from the vehicle’s surroundings, enabling autonomous driving and navigation.
However, the accurate perception and comprehension of the intricate dynamics within traffic environments pose several significant challenges: First, sensors may fail to provide data in certain situations, such as when cameras are obstructed. These data gaps can disrupt the continuous monitoring of traffic conditions, potentially leading to incomplete information. Second, sensor data collected in the traffic environment is subject to various sources of noise, including sensor inaccuracies, weather conditions, and lighting variations. This noise introduces uncertainty into the data, making it harder to derive precise information. Third, traffic environments are inherently complex and dynamic, involving numerous interacting factors such as vehicles, pedestrians, traffic signals, and road conditions. This complexity makes it challenging to accurately model and understand traffic behavior.
The advance of generative AI technology holds the potential to offer solutions to these challenges. We carried an in-depth literature review focusing on the application of generative AI to address challenges in traffic perception. The reviewed literature covers various aspects of traffic perception, including traffic data imputation, traffic estimation, traffic data analysis and traffic anomaly detection. Their relation is illustrated in Figure 3.

### v2

@@ 2.1.4. Energy-Based Models
EBMs (LeCun et al., 2006; Du and Mordatch, 2019) assign energy values to data points, with lower energy indicating a better fit. The goal is to assign lower energy to realistic points and higher energy to unrealistic ones, approximating the likelihood of observed data.
The probability density for data $\mathbf{x}$ is defined using a Boltzmann distribution: $p_{\theta}(\mathbf{x})=\frac{e^{-E_{\theta}(\boldsymbol{\mathbf{x}})}}{\int_{\tilde{\mathbf{x}}\in\mathcal{X}}e^{-E_{\theta}(\tilde{\mathbf{x}})}},$ where $E_{\theta}(\mathbf{x})$ is an energy function parameterized by $\theta$. To optimize EBMs, contrastive divergence compares positive samples (from real data) with negative samples (generated by the model), updating parameters to reduce energy for positives and increase energy for negatives. This is repeated iteratively to improve the model. Due to their flexibility and robustness, EBMs are used to model complex traffic patterns and perform multivariate time series forecasting in ITS applications (Yan et al., 2021).
@@ 2.1.5. Diffusion Probabilistic Model
DPMs (Sohl-Dickstein et al., 2015) generate data by reversing a noisy process to recover the original data step by step.
In the forward process, a Markov chain successively adds Gaussian noise to the data $\mathbf{x}_{0}$, resulting in an approximate posterior: $q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$
The reverse process starts from a noise distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$, and iteratively removes noise using learnable Gaussian transitions: $p_{\theta}(\mathbf{x}_{0:T})=p(\mathbf{x}_{T})\prod_{t=T}^{1}p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}).$
During training, the model predicts the noise added at each step, minimizing the difference between the predicted and actual noise. In the ITS domain, DPMs are utilized to model and generate complex traffic patterns by progressively denoising noisy traffic data. This enables the creation of realistic synthetic traffic scenarios, which are valuable for traffic simulation, prediction and decision-making.
@@ 2.1.6. Generative Models from Physical Process
GenPhys (Liu et al., 2023c) transforms physical partial differential equations (PDEs) into generative models, generalizing frameworks like diffusion models.
The goal is to generate new samples from observed data $p_{data}(\mathbf{x})$. A continuous physical process is described by: $\frac{d\mathbf{x}}{dt}=\mathbf{v}(\mathbf{x},t),$ which evolves the probability distribution $p(\mathbf{x},t)$ as: $\frac{\partial p(\mathbf{x},t)}{\partial t}+\nabla\cdot[p(\mathbf{x},t)\mathbf{v}(\mathbf{x},t)]-R(\mathbf{x},t)=0,$ where $R(\mathbf{x},t)$ represents particle birth or death. The goal is to design $p(\mathbf{x},t)$, $\mathbf{v}(\mathbf{x},t)$ such that $p(\mathbf{x},t)$ matches $p_{data}(\mathbf{x})$ at $t=0$ and becomes independent as $T\to\infty$.
To generate samples, the reverse physical process is used to evolve $\mathbf{x}(T)$, drawn from $p(\mathbf{x},T)$, backward to produce $\mathbf{x}(0)$. By incorporating GenPhys into ITS, we leverage the power of physics-informed generative modeling to improve traffic simulations, leading to better decision-making and more efficient transportation systems.
@@ 2.1.7. Generative Pre-trained Transformer
Generative Pre-trained Transformer (GPT) is a neural network designed for natural language processing (NLP) and text generation tasks. GPT uses a transformer decoder architecture (Vaswani et al., 2017) with self-attention and feedforward layers to capture long-range dependencies in text. It processes inputs with positional encodings and word embeddings, and generates text in an autoregressive manner, token by token. Pre-trained on large text datasets, GPT can be fine-tuned for specific tasks. These advanced capabilities can be leveraged to enhance various aspects of ITS, particularly in autonomous driving. For example, by fine-tuning GPT models on transportation-specific data, they can better understand complex driving scenarios, improving the decision-making processes of autonomous vehicles.
@@ 2.1.8. Comparative Analysis of Generative AI Techniques
In previous subsections, we introduced several major generative AI techniques, each with distinct advantages and limitations regarding performance, stability, flexibility, and computational requirements. For example, VAEs effectively learn latent representations suitable for data with inherent uncertainty, with moderate computational demands and stable training, but may struggle to generate high-quality images. GANs generate high-quality, diverse images using an adversarial process but face challenges like unstable training, mode collapse, and high computational demands. Table 2 summarizes these aspects for different generative AI techniques.
Technology Performance Stability Flexibility Computational Requirements
Variational Autoencoder (Kingma and Welling, 2013) Learns effective latent representations Suitable for data with inherent uncertainty May produce blurry images Relatively stable training process Capable of probabilistic generation Moderate computational resources needed
Generative Adversarial Network (Goodfellow et al., 2020) Generates high-quality and diverse images Training can be unstable Prone to mode collapse High flexibility in generation processes Requires substantial computational power
Normalizing Flow (Rezende and Mohamed, 2015) Accurately models complex probability distributions Stable but computationally intensive Supports both generation and inference High computational and parameter requirements for high-dimensional data
Energy-Based Model (LeCun et al., 2006) Models complex distributions effectively Robust to outliers Training can be slow and requires careful tuning Supports both generation and inference High computational demands during training
Diffusion Probabilistic Model (Sohl-Dickstein et al., 2015) Robust to noise and uncertainty Allows controlled generation Stable training process Flexible noise schedule Longer generation time and complex training
Generative Models from Physical Process (Liu et al., 2023c) Generates data grounded in physical processes Stability depends on the physical model accuracy Offers high interpretability Requires prior knowledge of physical parameters
Generative Pre-trained Transformer (Ghojogh and Ghodsi, 2020) Excels in natural language processing tasks Effective in modeling long-term dependencies Generally stable once trained Capable of generating continuous and coherent text Requires substantial computing resources for pre-training
Table 2. Comparison of Generative AI technologies evaluated in terms of performance, stability, flexibility, and computational requirements. Partly referenced by the work (Zhang et al., 2023d).
@@ 2.2. Important Generation Tasks
The advancement of generative technologies has provided impressive solutions for various generative tasks. In this section, we will explore the applications of generative technologies in ITS from the perspectives of text, images, videos, and cross-modal data, as shown in Figure 3.
Figure 3. The relation between different generation tasks in ITS.
@@ 2.2.1. Text Generation
Text generation focuses on using AI techniques to autonomously generate reasonable and meaningful text. In ITS, text generation is employed for tasks such as generating real-time traffic reports and automated incident alerts (Grigorev et al., 2024). This capability provides better insight into traffic conditions and enables timely responses to dynamic traffic situations, thereby improving overall traffic management and safety.
@@ 2.2.2. Image Generation
Image generation involves using AI algorithms to create new images or visual content based on given information. In the context of ITS, image generation is utilized for generating synthetic images for testing the driving algorithms in autonomous vehicles (Li et al., 2020b). By generating realistic images reflecting various traffic scenarios, weather conditions, and lighting environments, generative methods help improve the robustness and performance of driving behavior decision-making models used in ITS.
@@ 2.2.3. Video Generation
Video generation involves creating a series of frames that are semantically coherent to form a video. This is particularly important in ITS for simulating traffic scenarios, training autonomous driving systems, and analyzing vehicle behaviors. The ability to generate realistic traffic videos allows researchers and engineers to test and validate ITS algorithms under a wide range of conditions without the need for extensive real-world data collection (Zhao et al., 2024).
@@ 2.2.4. Cross-modality Data Generation
Cross-modality data generation refers to generating data in one modality based on features from a different modality. In ITS, this is crucial for tasks such as converting textual descriptions of traffic situations into visual representations (text-to-image), generating textual reports from traffic camera images (image-to-text), or creating simulations of traffic scenarios based on textual inputs (Ruan et al., 2024). These techniques enable more comprehensive traffic analysis and support the development of robust autonomous driving algorithms by providing diverse and rich datasets.
@@ 3. Generative AI for Traffic Perception
Traffic perception refers to the ability of ITS to collect and understand sensory information from the traffic environment. This information includes visual data from cameras, trajectory data from GPS-based devices, motion data from accelerometers, and environmental data such as weather and road conditions, as well as other traffic-related information like the presence of other vehicles, pedestrians, and road signs. Traffic perception is essential for making effective decisions and ensuring safety. In autonomous vehicles, traffic perception typically relies on cameras, LiDAR, millimeter-wave radar, and other technologies to collect and process data from the vehicle’s surroundings, enabling autonomous driving and navigation.
However, accurate perception and comprehension of the complex dynamics within traffic environments present several significant challenges. First, sensors may fail to provide data in certain situations, such as when cameras are obstructed. These data gaps can disrupt the continuous monitoring of traffic conditions, potentially leading to incomplete information. Second, sensor data collected in traffic environments is subject to various sources of noise, including sensor inaccuracies, weather conditions, and lighting variations. This noise introduces uncertainty, making it harder to derive precise information. Third, traffic environments are inherently complex and dynamic, involving numerous interacting factors such as vehicles, pedestrians, traffic signals, and road conditions, which accurately modeling and understanding traffic behavior challenging.
Recent advancements in generative AI technology offer promising solutions to these challenges. We carried an in-depth literature review focusing on the application of generative AI to address challenges in traffic perception. The reviewed literature covers various aspects of traffic perception, including data imputation, traffic estimation, data analysis and anomaly detection. Their relationships are illustrated in Figure 4.